### v1

# Multilinguality Does not Make Sense: Investigating Factors Behind Zero-Shot Transfer in Sense-Aware Tasks

Abstract
Cross-lingual transfer allows models to perform tasks in languages unseen during training and is often assumed to benefit from increased multilinguality. In this work, we challenge this assumption in the context of two underexplored, sense-aware tasks: polysemy disambiguation and lexical semantic change. Through a large-scale analysis across 28 languages, we show that multilingual training is neither necessary nor inherently beneficial for effective transfer. Instead, we find that confounding factors—such as fine-tuning data composition and evaluation artifacts—better account for the perceived advantages of multilinguality. Our findings call for more rigorous evaluations in multilingual NLP. We release fine-tuned models and benchmarks to support further research, with implications extending to low-resource and typologically diverse languages.
Multilinguality Does not Make Sense: Investigating Factors Behind Zero-Shot Transfer in Sense-Aware Tasks
Roksana Goworek1 Haim Dubossarsky1,2 1 Queen Mary University of London 2 Language Technology Lab, University of Cambridge {r.goworek, h.dubossarsky}@qmul.ac.uk
1 Introduction
Cross-lingual transfer enables models to leverage knowledge acquired in one language to perform tasks in another, and requires pre-training on multilingual corpora (e.g., Wu and Dredze, 2019; Ponti et al., 2018). This ability underpins many of today’s advances in multilingual NLP and has been evaluated across a wide range of tasks - from syntactic parsing and POS tagging to more complex, semantics-driven tasks like question answering, language inference, and paraphrasing (see Philippy et al. (2023) for a comprehensive review).
Polysemy disambiguation - despite being foundational to linguistic meaning and long-standing challenges in NLP (Navigli, 2009; Bevilacqua et al., 2021) - received little attention in transfer learning research. Polysemy poses unique challenges to transfer due to its inherently language-specific nature (Rzymski et al., 2020). For example, while the English word mole denotes ‘a small burrowing mammal’ and ‘a skin blemish,’ its Hindi counterpart
(til) refers to the latter sense but also means ‘sesame seed’. Such non-aligned senses across languages complicate direct transfer, which makes it an ideal testbed for evaluating the true limits of cross-lingual generalization.
A common assumption is that multilinguality itself, set by a model’s exposure to multiple languages, is key for cross-lingual transfer. While it is evident that models must support both source and target languages, the extent to which broader multilinguality facilitates transfer remains unclear and understudied. Does access to more languages intrinsically improve transfer, or do other factors Philippy et al. (2023), like language similarity, or training confounds, drive the observed gains?
In this work, we investigate the role of multilinguality in cross-lingual transfer for two underexplored, sense-aware tasks: polysemy disambiguation and lexical semantic change detection. We examine whether multilingual fine-tuning is truly essential for successful transfer, or whether previously reported benefits stem from confounding factors. Our results show that multilinguality is neither necessary nor intrinsically beneficial in these settings, challenging prevailing assumptions about the mechanisms underlying multilingual transfer.
While our findings may seem narrowly focused, they shed light on broader issues central to NLP, with implications for low-resource and domain-specific settings for which relying on broader multilingual models is not an option, as well as the wider research community. We release the best-performing models to support further analysis.
2 Related Work
Cross-lingual transfer studies have expanded from tasks like syntactic parsing, POS tagging, and semantic classification Wu and Dredze (2019); Choenni et al. (2023a); Pires et al. (2019); de Vries et al. (2022) to more complex ones such as NER, NLI Dolicki and Spanakis (2021); Srinivasan et al. (2021), and more recently QA, paraphrasing, and sentiment analysis Lauscher et al. (2020); Ahuja et al. (2023); Choenni et al. (2023b); Wang et al. (2023). Most studies rely on pretrained multilingual models like XLM-R and mBERT, typically fine-tuned on a source language and evaluated on target languages.
Research also focused on factors affecting transfer, including linguistic similarity Lauscher et al. (2020), pretraining corpus size and diversity Srinivasan et al. (2021); Ahuja et al. (2023), lexical overlap Patil et al. (2022); de Vries et al. (2022), and model architecture K et al. (2020). Language selection also varied, from high-resource Choenni et al. (2023b) to extremely low-resource languages tackled with continuous pretraining Ebrahimi and Kann (2021); Imani et al. (2023).
Despite extensive work, the role of multilinguality itself is rarely tested directly and is often assumed to be beneficial. A notable exception is Shaham et al. (2024), who compare PaLM 2 fine-tuned on monolingual vs. multilingual data for instruction tuning, and find modest multilingual exposure aids transfer, while too much can degrade it. However, their analysis is limited to a single model and task, leaving open questions about pretraining and generalizability. Choenni et al. (2023b) show source language examples can influence target predictions, but without explicitly controlling for multilinguality, offering only indirect evidence. Chang et al. (2024) directly manipulated multilinguality by pretraining over 10,000 models across 250 languages, but evaluated only with perplexity, omitting standard tasks and making comparisons difficult and replication impractical.
Notably underrepresented in transfer research are sense-aware tasks, such as polysemy disambiguation and lexical semantic change (LSC), despite the central role polysemy has in NLP. Unlike sentiment analysis or other tasks amenable to meaning-preserving translations, polysemy exhibits substantial language-specific variation Rzymski et al. (2020), making it particularly suitable for rigorous evaluation of zero-shot multilingual transfer ability.
Exceptions for polysemy are few, Raganato et al. (2020) reported transfer to 12 languages, but use only English as a source language, while Dairkee and Dubossarsky (2024) found near-zero transfer to Hindi, questioning the feasibility of zero-shot transfer in low-resource settings. Arefyev et al. (2021) showed that training on polysemy disambiguation generalizes to semantic change detection, linking the two tasks. Cassotti et al. (2023) fine-tuned models on multilingual polysemy datasets, achieving state-of-the-art results Schlechtweg et al. (2020) and strong transfer to unseen languages Periti and Tahmasebi (2024), though none of these works isolate the effect of multilinguality itself.
Overall, while multilingual transfer has been widely studied, its direct contribution - especially in lexically focused tasks like polysemy and LSC - remains underexplored. This work addresses that gap by systematically testing the role of multilinguality in sense-aware transfer. By manipulating multilingual conditions, we clarify when and how multilinguality supports cross-lingual generalization, helping to resolve prior conflicting findings and inform future research on transfer learning.
3 Methods
In this study we set to isolate multilinguality from confounds, enabling a clear assessment of its independent impact on zero-shot transfer in polysemy disambiguation and LSC tasks. We conduct a large-scale evaluation of zero-shot and full-shot cross-lingual transfer performance across 28 languages, focusing on the unique contribution of multilingual training. To achieve this, we implement an experimental framework that systematically controls for potential confounds, ensuring that observed effects are attributed to multilinguality rather than artifacts of training data size or pretraining exposure.
3.1 Tasks
The Word in Context (WiC) formulation of polysemy disambiguation is used. By pairing two sentences with the same polysemous word Pilehvar and Camacho-Collados (2019) transformed this task into a binary classification problem, where a target word appears either in the same sense or in a different sense as per the example below:
The couple went for a date last night.
He marked this date on my calendar.
This reformulation removed the need for a sense-label per word, which is language-specific, making the task more suitable for cross-lingual transfer.
For LSC, we compare the models’ representations of words occurring in natural sentences across two corpora from different time periods. The underlying assumption, which is the basis of all distributional semantics, is that changes in words’ meaning are reflected in measurable changes to their representations over time Periti and Montanelli (2024).
3.2 Data
MCL, XL and Hindi datasets, which together span 18 languages, were used for training and evaluation, while AM2iCO and LSCD, spanning 14 and 7 languages, respectively, were used only for evaluation.111AM2iCO uses English as a pivot language therefore is biased toward it; LSCD is not in the WiC format. As some languages have only development and test data, we followed Cassotti et al. (2023) who sampled 75% of the development data of each language to enable training on these languages (keeping the remaining 25% for setting hyperparameters), using his exact train-dev-test splits. All WiC datasets are class-balanced, setting the chance baseline at 50%. German, French and English are over represented in the dataset relative to other languages (see Figure 4).
MCL by Martelli et al. (2021) spans English, Arabic, French, Russian and Chinese, constructed by annotating sentences from native corpora: BabelNet Navigli and Ponzetto (2010), the United Nations Parallel Corpus Ziemski et al. (2016) and Wikipedia, with inter-annotator agreement of 0.95 and 0.9 on English and Russian, respectively.
XL by Raganato et al. (2020) used WordNet of Bulgarian, Chinese, Croatian, Danish, Dutch, English, Estonian, Japanese, Korean and Farsi, filtering out fine-grained senses. French, German and Italian used Wiktionary. Mean human accuracy was 80%, and varied between 74% for German, 87% for Danish and 97% for Farsi.
Hindi WiC by Dairkee and Dubossarsky (2024) consists of 12,000 sentence pairs, constructed from a sense-annotated Hindi Corpus Singh and Siddiqui (2016) of 60 polysemous nouns.
AM2iCO by Liu et al. (2021) has English paired with 14 target languages. Compiled from Wikipedia Dumps of each language by selecting words with at least two different Wikipedia pages that shows ambiguity in both the target language and English. Overall human accuracy was 90.6%, with an inter-annotator agreement of 88.4%.
Lexical Semantic Change Detection (LSCD) covers seven languages from different sources: English, German, Latin, and Swedish from Schlechtweg et al. (2020), Spanish from Zamora-Reina et al. (2022), Chinese from Chen et al. (2023), and Norwegian from Kutuzov et al. (2022).222Norwegian1, Norwegian2, are two corpora pairs, comparing the words from four time periods. For each language and target word, equal number of sentences were sampled from corpus 1 (historical) and corpus 2 (modern).
3.3 Multilingual Base Models
To ensure robustness, we use five multilingual models that differ in their language coverage, and in the pretraining proportions of these languages.
XLM-R Conneau et al. (2020) and mBERT Devlin et al. (2019) are pretrained on 100 and 104 languages, respectively, covering all languages in our study, which enables us to train and evaluate on all languages in zero- and full-shot transfer. Both models are commonly used in multilingual research, XLM-R was extensively used in the context of WiC and LSC (Raganato et al., 2020; Cassotti et al., 2023; Dairkee and Dubossarsky, 2024), providing a strong baseline for comparison.
BLOOM (BigScience Large Open-science Open-access Multilingual Language Model) by Le Scao et al. (2023) is pretrained on 46 languages, with a focus on low-resource languages, 10 of which overlap with those used in our study.
LLaMA3-8B-Instruct Grattafiori et al. (2024); AI@Meta (2024) is a state-of-the-art instruction-tuned language model with 8 billion parameters, significantly larger than the other models. Pretraining data is not public, and was assumed to include major NLP datasets Sainz et al. (2023), like WiC.
MuRIL (Multilingual Representations for Indian Languages) by Khanuja et al. (2021) is pretrained on 16 Indian languages + English, which, along with Hindi, are the only overlapping languages with our study. It was used to test an edge-case of zero-shot transfer (see §4.2).
3.4 Model Training and Testing
We fine-tune models on different combinations of the data, creating three conditions: (1) monolingual models, trained on a single language. Only the 5 largest languages were considered; (2) multilingual models, trained on all languages; and (3) multi⊗lingual models, trained on all but one held-out language. We also compare these models while controlling for the total amount of fine-tuning in the fixed f.t. condition, denoted by model=. By systematically comparing performance across conditions while controlling for fine-tuning size and pretraining exposure, we isolate the effect of multilinguality on cross-lingual transfer in WiC and LSC tasks.
Fixed Fine-Tuning
To control for the different training sizes across languages, fixed-size versions of datasets were created for the multi, english, french and german datasets by randomly subsampling 8,750 examples from their training sets.333Other languages were too small to subsample.
Finetuning
All models were finetuned in the same way, selectively focusing on different subsets of the WiC datasets.Following Cassotti et al. (2023) and after a failed pilot study with cross-encoders, a Siamese bi-encoder was used to generate two distinct vector representations (embeddings) for the two usages of the target word in the two sentences.444 Training hyperparameters and model sizes are reported in Appendix 6. The model outputs the cosine distance between the output embeddings of the two inputs, and, in order to adapt it to the binary nature of WiC, a threshold is applied to decide if the words are classified as having the same sense. The model is trained to update its parameters (i.e., embeddings) to maximize this distance when the target appears in different meanings (label 0) and minimize it when the meanings are the same in the two sentences (label 1), by minimizing contrastive loss. During training, as well as inference, special tokens, <t> and </t>, are placed around the target word in each sentence to signal what word the model should focus on. See Appendix I.2 for LLaMA’s finetuning details.
Testing
To evaluate models on WiC, a threshold for each model is determined by maximizing accuracy on the validation set of the training language. This threshold collapses the cosine distance between the output embeddings of the two inputs into a binary label. For LSC, we follow Cassotti et al. (2023) and output a fixed-size vector for the target word in each sentence. Taking all vectors for each word across the two time-bins, a change score is computed using APD and PRT measures Kutuzov and Giulianelli (2020), and evaluated against gold-label change scores using Spearman correlation, as standard in LSC Periti and Tahmasebi (2024).
4 Experiments and Evaluation
4.1 Multilinguality Effects and Confounds
To rigorously assess the role of multilinguality in zero-shot transfer, we conduct three types of controlled comparisons, ensuring that fine-tuning size is held constant and distinguishing between full-shot and true zero-shot transfer scenarios. We provide a comprehensive evaluation of four multilingual base-models (§3.3) on 28 languages, first examining whether multilingual training offers an advantage and then systematically accounting for key confounding factors. We summarize the average accuracies of the models in Figure 1 and present the true zero-shot transfer comparison in Table 2. The full results are reported in Appendix G. Due to space constraints, except for Figure 1, we show and discuss XLM-R results in the main text, only briefly discussing other models while providing their full results in the Appendices.
Testing for Multilingual Advantage
We contrast between mono models, which are fine-tuned on a single language, and the multi model, trained on all available data in the MCL and XL WiC datasets. If multilinguality is indeed critical, then multi should consistently outperform mono models, which by definition do not have access to information outside of their own language beyond their pretraining stage. Our results show a mixed pattern, with many mono models on-par or even better than multi on the MCL, AM2iCO and HINDI datasets, and a multilinguality advantage on the XL dataset (blue bars of the Full F.T condition in Figure 1). LLaMA is the exception to the rule, though its results are 5%-10% points lower than XLM-R. We attribute the low transfer of Hindi and Chinese mono models to their smaller dataset sizes, and subsequently tested whether dataset size is a potential confound in our analyses.
Dataset Size Confound
multi is fine-tuned on an order of magnitude more data than any other model (see Figure 4), which gives it an unfair advantage over all monolingual models. We therefore fix the size of the training datasets across all models (§3.4) to enable a fair comparison, and repeat the same analysis under these controlled conditions. The results in Figure 1, and especially the comparison between Full F.T and Fixed F.T (blue and orange bars, respectively) which are summarized in Table 1, show that the performance of the multi model drops much more than mono English when training size is controlled, perhaps due to a larger relative drop in training size, making mono English the best model across most datasets. Similar drops are observed for mBERT, BLOOM, and LLaMA (Tables 7-9), which for the latter largely diminished the advantage it had in Full F.T . Despite its drop in the Fixed F.T condition, LLaMA seems to benefit from multilingual fine-tuning more than other models. We attribute this to its nature as a generative model and pretraining uncertainty which may include data contamination unwanted for our experiments Sainz et al. (2023). Additionally, LLaMA relies on English prompts with explicit task instructions, which showed improved performance, particularly for hindi and chinese, where training data is more limited. Overall, this contrast (blue and orange) undermine the notion that training on a multilingual dataset inherently improves transfer. Instead, the advantage originally observed for multi in the Full F.T may largely be due to more training data (122.4k examples for multi cf. 54.7k for German, 46.1k for French, and 15.1k for English; see Table 4). This finding does not imply that training on more examples is not a good strategy to improve transfer, which it clearly is, only that attributing the performance gains to multilinguality is flawed.
Error Analysis
If multilinguality provides a novel sense-understanding, then not only should its performance improve, but also its errors should differ from monolingual models that lack exposure to such linguistic diversity. Instead, Figure 6 shows greater overlap of errors between multi and mono models, which even increases for Fixed F.T, supporting a notion that multi is simply another mono model (see formulation of alignment measures in Appendix E). This provides converging evidence that supports previous findings, further undermining the importance of multilinguality.
| MCL | XL | Hindi | Am2iCO | |
| multi | -5.0 | -5.0 | -2.7 | -4.0 |
| english | -0.6 | -0.6 | 1.2 | -1.5 |
| german | -6.8 | -4.4 | -2.0 | -4.3 |
| french | -9.5 | -5.5 | -19.6 | -8.7 |
| Language |
German |
French |
French |
English |
English |
Hindi |
Chinese |
Chinese |
| Dataset |
XL |
MCL |
XL |
MCL |
XL |
Hindi |
MCL |
XL |
| z.s. multi⊗= | 67.6 | 79.0 | 65.3 | 82.9 | 68.7 | 76.6 | 78.3 | 72.2 |
| german= | 74.1 | 77.0 | 63.8 | 78.3 | 60.6 | 74.7 | 75.0 | 67.5 |
| french= | 68.6 | 73.7 | 65.5 | 77.0 | 63.0 | 59.3 | 71.2 | 66.9 |
| english= | 69.3 | 84.1 | 65.6 | 89.2 | 68.9 | 79.9 | 78.4 | 76.2 |
| hindi= | 62.1 | 53.8 | 59.7 | 66.0 | 56.8 | 88.9 | 60.7 | 66.4 |
| chinese= | 62.3 | 61.3 | 58.1 | 68.1 | 61.3 | 44.2 | 70.8 | 69.4 |
Full-shot Exposure Artifact
We currently measure transfer by conflating zero- and full-shot, when the target language is present in training, together. Since multi is trained on all the languages it is later evaluated on (except Hindi), it is effectively evaluated under full-shot conditions. In contrast, mono models are only full-shot with respect to their training language, and zero-shot for all the rest. Thus, comparing transfer between multi and mono models is unfair, as full-shot learning is expected to be much better than zero-shot transfer.
To address this, we train five different multi⊗ models, excluding one language at a time on which that specific model is later evaluated on, while still fixing the amount of fine-tuning data as before, and compare them to mono models. We focus on the 5 languages with the most training data. If multilinguality is truly of merit, then multi models, that were trained on 14 languages overall (minus the held-out language they are evaluated on), should outperform mono models.
Table 2 shows that mono English outperforms zero-shot multi⊗ on all languages, and the only case where zero-shot multi⊗ outperforms mono models is English, where the english model is full-shot, and thus not considered in this evaluation (grayed). This further disproves the assumed benefit of multilingual training, as prior advantage in performance that was originally associated with multilinguality stemmed, at least in part, from mixing full-shot with zero-shot transfer conditions, which only multi possessed. Similar results were obtained for mBERT, BLOOM and even LLaMA, although with german as the best zero-shot model (see Tables 10-12). Interestingly, the strong zero-shot performance of multi models on English could be related to the prevalence of English in the base models’ pretraining data.
4.2 Underlying Factors of Successful Transfer
Correlation with Model’s Pretraining Size
Our analyses reject multilinguality as a key driver of zero-shot transfer, showing its effects stem from training data size - a confound, albeit a beneficial one. Yet, even after controlling for this, transfer results varied considerably across target languages.
Before their fine-tuning, multilingual models undergo pretraining on large-scale datasets spanning multiple languages. The prevalence of a target language within a model’s pretraining corpus, or "pretraining size", may influence the model’s ability to represent, process and transfer to that language.
To test this hypothesis, and understand what drives variation in transfer results, we computed Pearson correlations between languages’ log-transformed pretraining sizes in models for which pretraining sizes were available (XLM-R, mBERT, BLOOM) and their corresponding accuracies on the four WiC datasets and on LSCD.555Pretraining sizes are provided in Appendix 13. For BLOOM, we removed languages not in its pretraining.
Figure 2 shows strong correlations between the pretraining size of a language and its accuracy as a target language. We attribute BLOOM’s poor performance to its unusual pretraining language distribution that focuses on low-resource languages. The lack of correlation in XL remains unclear and requires further investigation (see §5 for discussion).
Linguistic Similarity
We tested the link between linguistic similarity and transferability using syntactic similarity scores Littell et al. (2017), which quantify similarity based on shared syntactic features, as is standard in NLP Philippy et al. (2023). Pearson correlations between similarity scores and model accuracies were computed across target language pairs. Syntactic similarity showed some correlation with zero-shot performance, the effect was weaker and less consistent than that of pretraining size (see Appendix §5). Moreover, we found that syntactic similarity is highly correlated with pretraining size, further undermining its contribution.
Zero-shot Transfer from Unknown Languages
While the correlation analysis provides useful insights, it remains slightly inconclusive. A key limitation is that we cannot pretrain models and directly manipulate this variable for a controlled experiment. We also lack sufficient fine-tuning data for languages that were less prominent during pretraining of our models, making it difficult to determine whether poor model performance stems from limited fine-tuning or limited exposure during pretraining.
To counter this, we analyzed transfer in an edge case scenario. We used MuRIL (§3.3), a model pretrained exclusively on 16 Indian languages and English. Evaluating zero-shot transfer performance of models that are fine-tuned on languages entirely absent from MuRIL’s pretraining corpus, can offer insight on the role of pretraining size in transfer.
Figure 3 shows graded transfer effects associated with the pretraining size, supporting the correlation analysis above. Hindi, being the dominant language in MuRIL (despite less data than English, it benefits from exposure to 16 Indic languages, including 2–4 with the same script and 9 with lexical overlap), enjoys the largest zero-shot transfer regardless of the fine-tuning language (tallest blue bars across all different models, not considering the full-shot condition in white), followed by English (yellow bars). Second, we observe substantial transfer from French, German and even Chinese on Hindi, languages MuRIL was not pretrained on (see §5 for further discussion). Third, MuRIL was able to learn and perform well on German, despite being absent during its pretraining stage.
We find that the multi model, despite being trained predominantly on German, French, English, and Chinese, failed to learn any meaningful sense disambiguation, let alone transfer them to Hindi. This highlights a stark contrast: while MuRIL trained on monolingual data can effectively learn the task from languages it was not pretrained on, the multilingual data appears too noisy for effective transfer in this extreme zero-shot setting.
4.3 Lexical Semantic Change Detection
| LSCD Langusage |
xl-lexeme |
multi |
german |
french |
english |
hindi |
chinese |
| English | .757 | .703 | .737 | .681 | .772 | .436 | .673 |
| German | .873 | .863 | .841 | .867 | .844 | .635 | .641 |
| Swedish | .754 | .801 | .754 | .618 | .724 | .480 | .489 |
| Latin | -.035 | .117 | .161 | .136 | .135 | -.177 | .091 |
| Spanish | .665 | .696 | .670 | .664 | .711 | .354 | .383 |
| Chinese | .734 | .652 | .649 | .499 | .737 | .524 | .593 |
| Norwegian1 | .668 | .729 | .638 | .697 | .777 | .525 | .400 |
| Norwegian2 | .634 | .655 | .604 | .580 | .645 | .433 | .439 |
| Average | .631 | .652 | .632 | .593 | .668 | .401 | .464 |
We evaluate XL-LEXEME, trained on multilingual and interlingual data (all WiC datasets plus AM2iCO), against the same mono models from earlier analyses. As in WiC, multi is matched or outperformed in all but two languages by mono English or German. XL-LEXEME outperforms only on German, but all well-trained mono models perform well on this task. Results correlate strongly with pretraining sizes in all three models (Figure 2), echoing WiC patterns.
Zero-shot transfer performs on par with full-shot. On German, mono French and English outperform the German model, and English is most effective on Spanish, Chinese, and Norwegian. PRT confirms these trends, with English showing the highest correlation. Across mBERT and BLOOM, mono models often match or outperform multi models (Tables 24-26).666LLaMA models were not evaluated on LSCD as prompting generative models with hundreds of usages to produce a scalar ‘change score’ is not practically feasible.
Overall, our results show that the best results in LSC are obtained by monolingual models, dismissing multilinguality as an important factor in cross-lingual transfer also on this task.
5 Discussion
Quantity, not Multilinguality
The multi model only partly outperformed monolingual models, with the english model often matching or surpassing it, even under the full f.t. condition. Strict control conditions across tests show that multilinguality does not inherently improve zero-shot transfer. Instead, its performance boost is due to factors like fine-tuning size and full-shot conditions. These results challenge the assumption that training on multilingual data inherently aids transfer; it is the quantity of training examples, not their multilinguality, that improves performance.
Quality, not only Quantity
mono English frequently outperformed multi model on most WiC and LSCD tasks, suggesting a special advantage. These results hold not only for fixed f.t, where quantity is equal, but even for full f.t, where english has far less training data compared to other languages. We suggest this advantage reflects the higher quality of English WiC, likely due to differences in dataset curation, annotation protocols, sense granularity, and text quality, echoing the conclusions from Philippy et al. (2023)’s survey.
Underlying Factors of Successful Transfer
The strong performance on English tasks observed in our results is likely due to its prevalence during pretraining, and supported by the strong correlations with PT-size reported in §4.2. The contrasting results from MuRIL, where transfer was strongest to Hindi – when it had the PT-size advantage over English – further underscore this point, providing an informative exception.
The influence of pretraining corpus size on cross-lingual transfer has been previously established Lauscher et al. (2020); Srinivasan et al. (2021); Ahuja et al. (2023), as well as the correlation between language similarity and transfer performance Wu and Dredze (2019); Pires et al. (2019); K et al. (2020). However, these factors affect tasks differently: pretraining size correlates more with semantic tasks (e.g., NLI, QA), while linguistic similarity benefits syntactic tasks (e.g., POS tagging, dependency parsing). This contrast is supported by results from Lauscher et al. (2020) and echoes Philippy et al. (2023) who argued that transfer is shaped by a combination of interacting factors.
This distinction sheds light on the types of linguistic knowledge that facilitate transfer. We propose that syntactic tasks rely more heavily on structural similarities across languages, and thus benefit from linguistic proximity. As syntactic patterns diverge across typologically distant languages, their utility in transfer diminishes. In contrast, semantic tasks - particularly those involving complex reasoning - are more dependent on the diversity and scale of pretraining data, which exposes models to a wider range of meaning representations. We suggest that sense-aware tasks, such as polysemy disambiguation and semantic change, follow this semantic pattern. Given the language-specific nature of word senses, syntactic similarity offers limited benefit of knowledge to transfer. Our findings confirm this: successful transfer in these tasks depends more on the size of multilingual pretraining than on linguistic similarity.
When More is Less
We include LLaMA in our analysis as a representative of modern generative large language models. It has more than 14 times more parameters than XLM-R, and was pretrained on orders of magnitude more data, likely including data very similar to our own Sainz et al. (2023). Nevertheless, LLaMA consistently underperforms relative to XLM-R across the majority of languages. This further supports our conclusion that training on more data and languages, even with the most modern models, does not guarantee improved cross-lingual transfer.
6 Conclusion
This study is the first to directly investigate the role of multilinguality in transfer for sense-aware tasks, polysemy and lexical semantic change. Our results indicate that the improved performance typically attributed to multilinguality largely stems from confounding factors such as training size, rather than linguistic diversity. Indeed, training on all available languages increases data quantity, inadvertently benefiting transfer, but this effect should not be mistakenly credited to multilingual diversity itself.
While further research is needed to determine whether these patterns generalize across other tasks, they may underline broader implications for multilingual NLP research that traditionally favored multilingual quantity over data quality. As expanding datasets across multiple languages is costly, and sometimes not practical for some tasks and languages, future research should prioritize understanding the characteristics of high-quality datasets, exposing potential confounds, and optimizing training resources for effective cross-lingual transfer.
7 Limitations
Zero-shot transfer enables models to perform tasks in languages without task-specific training data. However, its effectiveness depends on the pretrained model’s ability to capture the linguistic properties of both the source and target languages, as well as the availability of sufficient, high-quality data in the fine-tuning language. Our study is constrained by this limitation, as we were only able to train models on a limited set of languages with highly imbalanced data distributions. As in broader low-resource language research, high-resource languages dominate, while lower-resource languages often lack sufficient data for effective fine-tuning.
We evaluated zero-shot transfer on two tasks: WiC and LSCD. Both use the same bi-encoder architecture and cosine similarity to produce their outputs. While these results provide useful insights, further research is needed to validate the findings across other tasks and model architectures.
Our evaluation of LLaMA is limited to the WiC task, and several factors constrain its interpretability. First, its pretraining data is unknown so we could not assess the correlation of performance with pretraining size. Second, LLaMA’s pretraining could include the datasets used in our training or evaluation. And lastly, adapting WiC to LLaMA’s generative format (see Appendix I.2) introduced further limitations, such as the lack of a threshold to control sense discrimination, and a cue about the task for LLaMA, which encoder-based models lacked. Considering these, LLaMA is not directly comparable to the other models. Future work is needed to better understand these architectural differences.
Despite the exclusion of most LLMs, our study remains broadly relevant. Simpler encoder-based models continue to play a key role, especially in scenarios requiring efficient, resource-conscious deployment. Unlike LLMs, they can run on edge devices, perform better in low-resource settings and specialized domains, and are far more cost-effective. As LLMs remain computationally expensive and inaccessible in many real-world applications, encoder models will continue to be important for both research and industry.
Further investigation into the factors influencing zero-shot performance, particularly in low-resource languages, is essential for improving cross-lingual transfer and democratizing access to language technologies.
8 Ethical Considerations
This study investigates factors influencing successful zero-shot transfer, with the goal of informing the NLP community on how to develop more effective methods for speakers of low-resource languages. We evaluate models on tasks in a total of 28 languages and provide details on the languages and dataset sources used to create the datasets.
We do not foresee any direct ethical risks arising from our findings. Rather, this work promotes more responsible resource allocation by encouraging a shift away from continual dataset creation in favor of improving cross-lingual transfer techniques and understanding what defines high-quality datasets. However, it is important to acknowledge that zero-shot methods, while beneficial, may still introduce biases due to disparities in pretraining data, potentially disadvantaging underrepresented languages. Additionally, reliance on transfer from high-resource languages may reinforce linguistic hierarchies, where certain languages disproportionately influence model behavior.
Future work should continue to critically assess the impact of cross-lingual transfer on linguistic diversity and ensure that improvements in NLP benefit a wide range of language communities equitably.
9 Acknowledgments
We are grateful to Francesco Periti and Dominik Schlechtweg for their invaluable feedback and insightful advice on this paper.
This work has in part been funded by the research program Change is Key! supported by Riksbankens Jubileumsfond (under reference number M21-0021).
Appendix A Model Training Details
| Model Type | Model Name | Training Size |
| multi | multi | 122.4k |
| mono | german | 54.7k |
| french | 46.1k | |
| english | 15.1k | |
| hindi | 7k | |
| chinese | 2.5k |
| Model | Number of Parameters |
| XLM-R | 560M |
| BLOOM | 560M |
| mBERT | 178M |
| LLaMA3-8B-Instruct | 8B |
| MuRIL | 506M |
| Hyperparameter | Value |
| Hidden activation | gelu |
| Hidden dropout probability | 0.1 |
| Hidden size | 1024 |
| Initializer range | 0.02 |
| Intermediate size | 4096 |
| Layer norm epsilon | |
| Max position embeddings | 514 |
| Number of attention heads | 16 |
| Number of hidden layers | 24 |
| Position embedding type | Absolute |
| Vocabulary size | 250004 |
| Learning rate | |
| Weight decay | 0.0 |
| Max sequence length () | 128 |
Appendix B Percentage Change Between Full F.T. - Fixed F.T.
| MCL | XL | Hindi | Am2iCO | |
| multi | -3.3 | -5.1 | -12.3 | -2.4 |
| english | 1.6 | 2.1 | 0.1 | 1.5 |
| german | 1.6 | 1.5 | -18.2 | -3.0 |
| french | -7.7 | 0.3 | 23.1 | -2.8 |
| MCL | XL | Hindi | Am2iCO | |
| multi | -11.5 | -9.1 | -13.0 | -4.8 |
| english | -5.3 | -0.9 | -24.9 | -3.2 |
| german | -11.5 | -5.8 | -20.4 | -3.7 |
| french | -17.6 | -5.4 | -24.5 | -4.0 |
| MCL | XL | Hindi | Am2iCO | |
| multi | -5.9 | -2.2 | 12.8 | 0.5 |
| english | -1.9 | 3.5 | 3.6 | -2.1 |
| german | -1.4 | -1.5 | -18.3 | 0.7 |
| french | -1.1 | -0.8 | 4.8 | -2.6 |
Appendix C Linguistic Similarity
Appendix D Zero-Shot Comparison of mBERT and BLOOM Models
| Language |
German |
French |
French |
English |
English |
Hindi |
Chinese |
Chinese |
| Dataset |
XL |
MCL |
XL |
MCL |
XL |
Hindi |
MCL |
XL |
| z.s. multi⊗= | 62.9 | 74.9 | 62.7 | 69.9 | 64.6 | 58.4 | 65.8 | 64.5 |
| german= | 73.9 | 72.2 | 61.3 | 73.6 | 56.3 | 51.0 | 67.7 | 63.9 |
| french= | 66.2 | 66.0 | 65.9 | 72.5 | 62.8 | 46.2 | 58.2 | 60.4 |
| english= | 65.6 | 78.3 | 62.8 | 84.3 | 63.9 | 73.4 | 73.0 | 68.1 |
| hindi= | 55.7 | 67.0 | 56.4 | 69.6 | 55.9 | 76.1 | 64.1 | 60.7 |
| chinese= | 61.9 | 72.7 | 63.4 | 78.1 | 61.4 | 59.7 | 70.8 | 68.0 |
| Language |
German |
French |
French |
English |
English |
Hindi |
Chinese |
Chinese |
| Dataset |
XL |
MCL |
XL |
MCL |
XL |
Hindi |
MCL |
XL |
| z.s. multi⊗= | 48.6 | 54.1 | 52.9 | 54.1 | 53.9 | 57.9 | 50.0 | 50.1 |
| german= | 61.3 | 49.4 | 50.7 | 49.7 | 51.0 | 48.4 | 50.0 | 50.0 |
| french= | 50.3 | 52.3 | 60.4 | 51.1 | 53.1 | 45.1 | 50.0 | 50.0 |
| english= | 51.2 | 67.4 | 57.1 | 74.6 | 58.6 | 48.2 | 55.6 | 52.8 |
| hindi= | 49.7 | 50.0 | 50.1 | 50.0 | 49.9 | 88.6 | 50.0 | 50.0 |
| chinese= | 49.8 | 50.0 | 50.1 | 50.0 | 50.0 | 46.2 | 54.5 | 59 |
| Language |
German |
French |
French |
English |
English |
Hindi |
Chinese |
Chinese |
| Dataset |
XL |
MCL |
XL |
MCL |
XL |
Hindi |
MCL |
XL |
| z.s. multi⊗= | 66.7 | 77.6 | 67.4 | 74.1 | 68.4 | 71.5 | 72.7 | 68.9 |
| german= | 76.6 | 77.1 | 68.0 | 77.5 | 64.9 | 47.2 | 75.9 | 71.9 |
| french= | 68.6 | 70.5 | 65.8 | 77.1 | 67.8 | 64.5 | 74.5 | 71.6 |
| english= | 63.6 | 74.2 | 64.5 | 84.0 | 70.2 | 66.2 | 71.1 | 64.4 |
| hindi= | 59.8 | 66.4 | 56.8 | 72.7 | 64.1 | 72.9 | 69.4 | 64.7 |
| chinese= | 64.0 | 70.9 | 62.1 | 70.7 | 68.7 | 68.5 | 71.5 | 70.2 |
Appendix E Error Analysis
We perform error analysis to further analyse how, and if multilingual models differ from their monolingual counterparts.
Let:
be the set of monolingual models in the same fine-tuning condition,
be the model of interest, where ,
be the set of incorrect predictions made by model ,
, the proportion of ’s errors also made by model .
Then, the average proportion of alignment of model with the monolingual models is given by:
Appendix F Pretraining Sizes of Languages Used in Our Analysis
| ISO Code | Language | XLM-R | mBERT | BLOOM | MuRIL |
| AR | Arabic | 3.37 | 0.72 | 4.26 | - |
| BG | Bulgarian | 4.07 | 0.24 | - | - |
| BN | Bengali | 2.24 | 0.12 | 2.91 | - |
| DA | Danish | 3.84 | 0.24 | - | - |
| DE | German | 4.21 | 1.66 | - | - |
| EN | English | 5.71 | 2.89 | 6.12 | 3.30 |
| ES | Spanish | 3.99 | 1.66 | 5.10 | - |
| ET | Estonian | 1.96 | 0.24 | - | - |
| EU | Basque | 1.10 | 0.12 | 1.16 | - |
| FA | Persian | 4.72 | 0.43 | - | - |
| FI | Finnish | 4.01 | 0.43 | - | - |
| FR | French | 4.06 | 1.66 | 5.27 | - |
| HI | Hindi | 3.05 | 0.12 | 3.18 | 1.95 |
| HR | Croatian | 3.07 | 0.24 | - | - |
| ID | Indonesian | 5.01 | 0.43 | 2.98 | - |
| IT | Italian | 3.44 | 1.14 | - | - |
| JA | Japanese | 4.25 | 1.14 | - | - |
| KA | Georgian | 2.31 | 0.12 | - | - |
| KK | Kazakh | 2.00 | 0.12 | - | - |
| KO | Korean | 4.01 | 0.43 | - | - |
| LA | Latin | 1.25 | 0.06 | - | - |
| NL | Dutch | 3.41 | 0.72 | - | - |
| NO | Norwegian | 3.91 | 0.43 | - | - |
| RU | Russian | 5.63 | 1.66 | - | - |
| SV | Swedish | 2.57 | 0.72 | - | - |
| TR | Turkish | 3.09 | 0.43 | - | - |
| UR | Urdu | 1.90 | 0.12 | 1.28 | 1.00 |
| ZH | Simplified Chinese | 3.87 | 1.14 | 5.50 | - |
| ZH | Traditional Chinese | 2.87 | 1.14 | 0.54 | - |
Language proportions for LLaMA-3–8B-Instruct’s pretraining corpus are not publicly available.
Appendix G All Results on Monolingual and Cross-Lingual WiC tasks
Type Dataset MCL XL Hindi AR EN FR RU ZH BG DA DE EN ET FA FR HR IT JA KO NL ZH HI multi multi 82.6 88.0 82.6 84.3 85.5 83.0 76.0 83.7 70.4 62.3 66.6 78.0 70.6 72.5 66.7 78.4 77.1 79.4 79.3 mono german 82.4 85.2 81.0 82.9 82.1 69.0 72.1 84.4 67.6 62.3 62.0 67.3 65.2 65.7 62.1 68.2 73.0 73.4 76.7 french 80.8 85.5 81.3 81.4 81.3 67.5 73.9 71.0 67.9 60.0 67.9 78.1 65.7 69.3 63.0 66.7 76.3 73.8 78.8 english 82.4 90.2 84.6 85.1 84.4 70.8 73.8 67.8 70.6 60.0 62.8 65.9 70.8 64.2 62.9 73.5 74.5 76.2 78.7 hindi 56.1 66.0 53.8 62.5 60.7 61.7 61.0 62.1 56.8 50.3 53.8 59.7 48.5 52.2 51.9 58.2 61.2 66.4 88.9 chinese 62.7 68.1 61.3 61.9 72.6 57.3 62.4 62.3 61.3 55.1 72.4 58.1 58.8 59.8 57.0 58.2 63.4 68.8 44.2 fixed fine-tuning multi= 77.6 83.1 80.7 77.6 79.0 69.0 70.3 75.1 69.0 63.1 73.8 67.2 69.6 65.2 65.5 70.2 68.7 73.2 76.6 german= 75.9 78.3 77 73.2 75 67.1 65.8 74.1 60.6 57.4 60.6 63.8 64.5 63 61.5 61.8 66.9 67.5 74.7 french= 70.5 77 73.7 70.6 71.2 59 64.9 68.6 63 56.9 72 65.5 63.5 63.5 61.5 59.2 64.6 66.9 59.3 english= 82.8 89.2 84.1 83.4 84.0 71.9 71.8 69.3 68.9 58.2 66.6 65.6 73.5 62.7 61.5 68.5 72.6 75.1 79.9 hindi 56.1 66.0 53.8 62.5 60.7 61.7 61.0 62.1 56.8 50.3 53.8 59.7 48.5 52.2 51.9 58.2 61.2 66.4 88.9 chinese 62.7 68.1 61.3 61.9 72.6 57.3 62.4 62.3 61.3 55.1 72.4 58.1 58.8 59.8 57.0 58.2 63.4 68.8 44.2 z.s. multi german⊗= 75.9 82.9 77.7 77.9 78.4 68.2 68.3 67.6 68.1 59 69.2 67.6 69.1 66 63.6 67.4 67.4 71.6 79 french⊗= 74.2 81.1 79 77.8 76.4 70.6 68.7 70.7 68.5 59.5 66.9 65.3 70.3 62.3 61.7 67.2 71 71.1 76.4 english⊗= 76.6 82.9 78.6 74.5 77.7 70.7 68 73.6 68.7 63.1 75.9 65.8 73 64.9 61.7 68.4 68 71.6 76.8 chinese⊗= 80.1 83.3 81.2 77.3 78.3 70.7 69.8 71.9 67.9 58.5 66.5 66.8 70.8 66.6 62.5 68.6 70.4 72.2 79.9
Type Dataset AM2iCO AR BN DE EU FI ID JA KA KK KO RU TR UR ZH multi multi 68.0 65.7 69.3 65.4 70.0 67.8 65.8 67.2 63.2 66.9 69.9 67.4 64.5 66.6 mono german 64.5 63.9 67.8 66.6 68.3 67.1 64.4 64.2 62.5 66.9 67.7 66.4 61.2 65.9 french 67.7 68.0 70.0 66.7 72.5 68.2 68.3 68.9 61.8 67.9 72.4 68.2 61.5 66.5 english 67.4 66.1 69.9 66.2 73.0 68.7 68.5 66.5 61.8 68.9 71.8 68.7 66.0 67.6 hindi 58.4 51.3 60.6 55.0 60.1 54.8 61.5 52.3 52.2 61.5 60.3 58.0 55.8 59.7 chinese 56.0 50.1 59.7 57.0 58.6 56.4 60.0 53.0 55.0 58.9 57.1 57.7 55.5 59.9 fixed fine-tuning mlcl= 61.0 62.7 66.3 62.1 67.2 64.1 64.8 61.0 58.2 63.2 65.9 62.4 60.5 61.8 german= 60.2 57.6 64.6 61.9 63.6 61.2 61.6 63.2 60.2 59.4 64.4 61.1 57.5 61.3 french= 60.7 52.6 63.4 57.4 63.4 60.0 62.1 57.0 56.2 59.8 62.0 59.4 57.0 56.1 english= 65.6 62.7 70.0 63.6 71.1 67.6 67.3 64.5 60.2 67.9 70.8 66.6 66.2 65.6 hindi 58.4 51.3 60.6 55.0 60.1 54.8 61.5 52.3 52.2 61.5 60.3 58.0 55.8 59.7 chinese 56.0 50.1 59.7 57.0 58.6 56.4 60.0 53.0 55.0 58.9 57.1 57.7 55.5 59.9 z.s. multi german⊗= 68.2 68.0 70.9 68.7 75.3 70.4 69.0 70.3 64.8 69.8 73.8 68.2 65.8 66.6 french⊗= 72.0 69.7 72.5 69.9 75.3 72.4 72.6 72.6 68.0 73.1 74.6 72.7 68.2 69.1 english⊗= 62.7 60.4 64.9 59.3 66.5 62.5 62.9 60.8 57.8 63.3 64.7 62.1 62.2 62.8 chinese⊗= 71.1 72.3 72.3 68.1 76.9 70.1 69.9 70.4 68.0 72.5 74.4 72.2 68.8 67.5
Type Dataset MCL XL Hindi AR EN FR RU ZH BG DA DE EN ET FA FR HR IT JA KO NL ZH HI multi multi 75.1 81.4 76.2 76.7 74.8 79.5 65.9 84.8 64.1 56.2 64.4 74.6 68.6 72.0 61.2 67.1 69.7 72.2 70.7 mono german 65.1 72.5 68.4 68.2 67.6 55.6 59.3 82.9 57.6 55.4 59.1 58.3 57.1 57.1 54.6 54.2 57.5 61.6 69.3 french 71.6 74.6 71.7 71.9 69.7 57.9 62.2 63.6 60.4 58.7 60.1 76.2 56.9 63.5 57.9 57.7 65.2 65.5 69.3 english 73.5 82.1 76.1 73.2 72.7 57.9 63.3 61.9 64.1 52.8 62.1 61.5 61.8 59.6 57.6 54.8 63.4 66.4 73.3 hindi 64.7 69.6 67 65.5 64.1 51.1 57.5 55.7 55.9 48.2 59.8 56.4 53.2 52.9 50.7 51 52 60.7 76.1 chinese 68.7 78.1 72.7 69.1 70.8 56.4 63 61.9 61.4 56.7 65.9 63.4 60.5 60 56.9 55.9 63.6 68 59.7 fixed fine-tuning multi= 71.5 79.3 76.1 70.6 70.2 63.9 62.1 72.4 64.2 60.3 66 65.4 66.7 63.7 60.9 58.8 63.6 65.7 58.4 german= 69.4 73.6 72.2 67.1 67.7 59.6 59.2 73.9 56.3 56.9 63.9 61.3 59.3 61.8 58.7 58.3 57.3 63.9 51.0 french= 61.4 72.5 66.0 62.9 58.2 61.2 62.8 66.2 62.8 57.2 67.5 65.9 64.5 62.8 55.7 58.3 64.1 60.4 46.2 english= 74.6 84.3 78.3 75.4 73 62.5 62.3 65.6 63.9 54.1 64.6 62.8 67.9 60 58.7 59.4 64.8 68.1 73.4 hindi 64.7 69.6 67 65.5 64.1 51.1 57.5 55.7 55.9 48.2 59.8 56.4 53.2 52.9 50.7 51 52 60.7 76.1 chinese 68.7 78.1 72.7 69.1 70.8 56.4 63 61.9 61.4 56.7 65.9 63.4 60.5 60 56.9 55.9 63.6 68 59.7 z.s. multi german⊗= 72.1 76.2 74.4 69.3 69.2 61.3 61.9 62.9 64.5 54.4 65.0 63.2 68.9 61.7 59.5 60.4 62.0 65.9 61.5 french⊗= 65.5 74.2 74.9 70.0 67.5 63.0 61.5 68.2 64.1 56.2 65.8 62.7 66.7 60.5 57.9 58.5 60.5 63.3 64.7 english⊗= 57.3 69.9 63.8 62.2 63.8 61.2 59.1 69.4 64.6 56.7 63.2 64.2 60.0 62.7 59.0 60.1 64.9 63.2 43.8 chinese⊗= 66.2 73.9 71.6 68.0 65.8 63.4 61.7 69.1 64.3 57.9 62.8 62.2 65.7 62.5 60.7 60.0 64.1 64.5 58.3
Type Dataset AM2iCO AR BN DE EU FI ID JA KA KK KO RU TR UR ZH multi multi 60.5 61.6 64.9 59.1 64.4 62.7 60.4 59.9 57.2 59.7 64.8 61.1 57.8 60.9 mono german 56.7 60.1 61.9 59.6 61.9 61.7 59.7 58.1 59.2 57.8 62.3 61.7 57.2 58.7 french 58.5 57 62.4 61 64.4 64.1 62.6 58 55 59.6 64.2 62.4 57.5 60.7 english 59.8 60 63.8 60.6 63.9 64.7 61.5 61.2 59 62.4 63.9 61.6 61.5 61.5 hindi 51.2 50.9 57.2 50.1 50.8 52.2 49.9 50.2 49.2 50.8 53.8 50.8 51 50 chinese 53 52.7 59.7 50.2 52.7 55.9 52.9 51 48.8 52.9 58.8 52.8 50.7 51.5 fixed fine-tuning mlcl= 58.3 60.4 64.7 56.5 61.1 61.0 59.0 55.5 53.0 58.1 61.7 58.7 58.2 55.6 german= 55.8 55.9 63.5 55.1 59.1 60.6 56.6 54.6 51.0 57.3 60.0 57.8 53.5 53.4 french= 58.9 57.7 61.7 56.6 60.4 59.0 57.8 54.5 53.5 58.5 60.3 58.6 56.5 56.9 english= 62.8 62.6 66.6 60.3 65.3 65.1 63.7 62 59.5 63.9 65.7 63.1 63 62.8 hindi 51.2 50.9 57.2 50.1 50.8 52.2 49.9 50.2 49.2 50.8 53.8 50.8 51 50 chinese 53 52.7 59.7 50.2 52.7 55.9 52.9 51 48.8 52.9 58.8 52.8 50.7 51.5 z.s. multi german⊗= 62.3 63.1 66.4 61.9 68.2 63.8 64.5 63.0 60.0 64.6 65.8 65.2 61.2 65.6 french⊗= 65.4 64.6 68.9 64.8 71.0 69.1 66.1 64.4 65.2 64.8 67.5 68.6 64.2 67.6 english⊗= 59.2 57.9 60.8 56.8 60.4 60.4 60.8 57.0 55.8 58.5 61.0 60.0 57.0 60.1 chinese⊗= 63.3 66.0 69.0 62.8 68.6 65.1 64.0 64.9 60.2 65.7 65.8 67.1 59.5 65.5
Type Dataset MCL XL Hindi AR EN FR RU ZH BG DA DE EN ET FA FR HR IT JA KO NL ZH HI multi multi 71.4 73.9 65.8 54.8 75.6 74.0 56.1 80.8 62.3 58.7 55.1 72.3 54.2 66.7 57.5 69.3 59.5 71.0 70.9 mono german 62.9 66.1 62.3 55.2 66.1 51.8 54.9 82.9 58.4 55.9 54 57.5 51 53.9 56.4 56.2 55.5 61.2 68.8 french 72.7 74.9 68.4 54.6 72.8 52.8 50.1 52.7 61 52.3 55.9 74.4 50 57.9 54.4 56.4 52.8 66.2 69.7 english 68.7 74.1 70.4 53.3 68.8 50.1 50.4 50.2 62.4 51 52.9 60.1 50.7 52 52.9 50.1 49.4 59.1 73.1 hindi 55 50 50 50 50 50 50 49.7 49.9 50 49.2 50.1 49.8 50 50 50.9 50 50 88.6 chinese 61.8 50 50 50 54.5 49.5 50 49.8 50 49.5 53 50.1 49 50.2 50 52 49.9 59 46.2 fixed fine-tuning multi= 64 59.2 54.9 50.7 55.2 53.3 50.7 61.7 53.9 56.2 52.2 57.3 53.4 58.6 53.5 56.8 56.3 55.7 57.9 german= 57.3 49.7 49.4 48.5 50.0 49.9 50.6 61.3 51.0 46.9 50.5 50.7 54.2 50.8 50.2 54.7 53.5 50.0 48.4 french= 52.2 51.1 52.3 50.0 50.0 49.7 49.6 50.3 53.1 49.0 52.2 60.4 48.3 53.7 50.1 50.5 49.7 50.0 45.1 english= 60.8 74.6 67.4 50.5 55.6 50.7 51.5 51.2 58.6 50.3 50 57.1 50 54.6 51.1 52.7 49.2 52.8 48.2 hindi 55 50 50 50 50 50 50 49.7 49.9 50 49.2 50.1 49.8 50 50 50.9 50 50 88.6 chinese 61.8 50 50 50 54.5 49.5 50 49.8 50 49.5 53 50.1 49 50.2 50 52 49.9 59 46.2 z.s. multi german⊗= 52.5 55.5 54.2 48.9 54.5 51.8 50.3 48.6 57.5 51.5 53.2 57.1 52.7 58.3 53.2 58.8 54.0 52.8 54.7 french⊗= 55.9 52.6 50.7 51.8 52.9 47.5 51.2 55.7 55.9 53.8 49.4 52.9 52.7 58.4 53.2 56.0 50.5 52.9 48.1 english⊗= 62.8 54.1 54.3 52.9 57.1 50.9 51.7 58.3 53.9 56.2 50.7 56.7 49.5 54.2 53.4 57.7 52.6 56.0 44.4 chinese⊗= 50.5 51.0 54.7 50.1 50.0 51.4 49.9 57.2 54.8 54.6 51.2 56.0 53.7 57.4 51.1 56.0 54.4 50.1 43.7
Type Dataset AM2iCO AR BN DE EU FI ID JA KA KK KO RU TR UR ZH multi multi 58.0 56.6 55.2 54.0 52.3 57.4 56.2 52.4 53.2 53.6 51.8 52.8 55.0 59.6 mono german 55.2 55.6 52 54 52.6 53.4 55.5 51.8 57.2 49.3 53 53.3 54 55 french 57.8 54.1 55.5 51.3 50.9 58.4 53.6 49.1 50.7 51.4 51.2 50.4 57.5 60.8 english 59.2 54.7 55.8 52.8 53.4 56.3 55.4 51.4 51.8 49.9 52.6 50.3 55.8 55.1 hindi 50 50 50 49 50 50.1 50.1 50.1 49.8 50.4 50.1 50 50 50.4 chinese 50 50 50 47.5 49.9 50 51.4 50.1 49.5 49.8 50.9 49.9 50 48.2 fixed fine-tuning multi= 49.9 50.0 51.0 50.1 50.3 51.0 51.1 50.1 49.8 49.9 49.8 48.2 50.0 49.4 german= 50.0 50.0 50.0 50.2 49.2 50.3 50.7 50.1 50.0 48.9 48.9 49.8 50.0 51.5 french= 50.0 50.0 50.6 49.2 49.4 50.0 49.6 50.0 50.2 48.5 49.2 49.9 50.0 49.5 english= 55.5 50 52.3 50.7 51.5 50.4 50 50.1 50 49.7 50.1 49.5 50 49.9 hindi 50 50 50 49 50 50.1 50.1 50.1 49.8 50.4 50.1 50 50 50.4 chinese 50 50 50 47.5 49.9 50 51.4 50.1 49.5 49.8 50.9 49.9 50 48.2 z.s. multi german⊗= 52.2 49.6 51.4 51.1 50.0 51.4 50.8 52.5 53.0 50.5 49.2 51.3 50.2 53.1 french⊗= 51.0 51.4 51.5 52.7 50.0 52.7 52.8 53.3 49.2 51.1 50.8 52.5 51.5 49.0 english⊗= 50.1 50.0 49.8 50.0 51.4 50.7 50.4 50.0 50.0 49.9 50.0 50.4 50.0 47.5 chinese⊗= 52.7 50.1 54.1 50.9 51.3 49.4 50.6 50.5 50.2 51.5 48.1 51.9 50.2 49.5
Type Dataset MCL XL Hindi AR EN FR RU ZH BG DA DE EN ET FA FR HR IT JA KO NL ZH HI multi multi 71.8 85.6 76.5 78.1 80.2 69.5 67.9 77.9 71.6 62.1 79.0 74.2 69.4 70.9 66.9 73.7 75.1 72.8 58.7 mono german 71.5 78.9 77.0 75.9 77.9 64.3 68.9 80.0 70.2 53.3 69.0 69.0 67.6 64.4 63.6 65.6 71.2 72.4 65.5 french 70.4 78.7 72.7 71.7 75.3 62.8 67.0 69.9 69.6 53.3 69.2 76.5 63.0 67.6 63.1 68.6 72.7 70.3 59.7 english 70.3 85.0 76.9 74.8 70.9 54.9 62.3 60.3 67.7 50.5 63.2 63.2 55.1 57.1 53.2 55.6 59.8 65.4 62.6 hindi 64.8 72.7 66.4 68.7 69.4 56.4 58.3 59.8 64.1 50.8 67.6 56.8 60.0 58.4 57.5 58.2 62.8 64.7 72.9 chinese 69.3 70.7 70.9 71.5 71.5 64.8 63.7 64.0 68.7 56.7 70.9 62.1 67.2 61.7 60.1 66.4 65.4 70.2 68.5 fixed fine-tuning multi= 73.4 79.8 74.9 73.2 73.2 66.5 65.2 75.6 71.0 62.1 80.1 70.6 70.6 67.4 64.0 69.0 68.8 71.5 71.5 german= 70.9 77.5 77.1 71.7 75.9 60.5 65.3 76.6 64.9 57.4 74.1 68.0 66.2 64.4 61.2 59.8 69.2 71.9 47.2 french= 70.7 77.1 70.5 69.8 74.5 65.4 63.4 68.6 67.8 60.5 76.9 65.8 64.5 63.3 63.6 68.4 63.7 71.6 64.5 english= 68.5 84.0 74.2 74.7 71.1 62.7 62.4 63.6 70.2 55.1 63.2 64.5 59.8 61.0 58.6 62.8 65.9 64.4 66.2 hindi 64.8 72.7 66.4 68.7 69.4 56.4 58.3 59.8 64.1 50.8 67.6 56.8 60.0 58.4 57.5 58.2 62.8 64.7 72.9 chinese 69.3 70.7 70.9 71.5 71.5 64.8 63.7 64.0 68.7 56.7 70.9 62.1 67.2 61.7 60.1 66.4 65.4 70.2 68.5 z.s. multi german⊗= 63.2 81.4 74.8 70.8 70.6 64.6 64.3 66.7 71.4 62.1 78.1 70.4 66.9 66.2 62.0 66.4 71.4 65.7 67.6 french⊗= 69.3 78.9 77.6 74.1 76.6 65.7 63.6 74.2 72.2 63.1 77.5 67.4 67.6 65.5 64.6 69.6 69.7 70.4 61.6 english⊗= 61.6 74.1 69.9 67.2 69.6 63.2 64.2 76.3 68.4 58.5 73.9 69.4 63.5 66.4 62.1 68.7 69.6 67.6 56.7 chinese⊗= 70.4 78.3 72.8 71.7 72.7 63.1 62.9 72.0 68.2 65.1 71.1 67.8 68.9 62.7 62.7 67.5 68.8 68.9 62.9
Type Dataset AM2iCO AR BN DE EU FI ID JA KA KK KO RU TR UR ZH multi multi 61.1 59.7 63.5 58.1 60.0 59.8 59.7 57.4 56.0 61.7 60.3 58.6 55.2 61.5 mono german 59.9 55.0 60.6 55.2 57.7 57.3 60.0 55.0 56.8 60.2 58.3 58.3 54.5 57.3 french 58.5 58.0 62.5 57.1 63.6 58.8 59.4 56.5 53.2 60.5 60.3 60.2 57.2 59.0 english 61.9 61.0 63.9 57.6 63.9 60.1 62.0 59.5 57.0 63.6 62.3 60.1 59.5 62.1 hindi 59.8 59.7 62.1 56.1 60.0 59.8 59.1 55.8 55.2 59.5 60.5 57.0 55.5 58.6 chinese 59.0 56.4 62.2 55.4 59.4 60.1 58.8 53.2 53.5 56.3 59.8 59.3 55.5 58.5 fixed fine-tuning multi= 61.6 58.9 64.6 58.5 63.8 61.1 60.2 58.2 55.0 59.0 59.5 60.0 58.0 61.3 german= 57.1 57.7 60.3 56.7 60.5 59.3 59.5 58.5 55.2 60.8 58.9 57.5 55.8 58.0 french= 56.3 57.6 58.3 57.5 59.4 58.6 55.9 52.9 52.0 55.3 56.9 56.4 56.0 55.4 english= 60.1 59.1 61.1 56.5 60.7 58.6 58.1 60.3 56.0 59.2 60.4 56.9 58.5 59.6 hindi 59.8 59.7 62.1 56.1 60.0 59.8 59.1 55.8 55.2 59.5 60.5 57.0 55.5 58.6 chinese 59.0 56.4 62.2 55.4 59.4 60.1 58.8 53.2 53.5 56.3 59.8 59.3 55.5 58.5 z.s. multi german⊗= 73.1 73.0 78.9 72.1 78.6 76.3 75.8 72.4 71.5 76.0 80.0 75.8 69.8 76.3 french⊗= 71.3 68.7 72.5 69.6 73.8 71.1 70.1 69.7 68.5 72.2 74.1 71.6 67.5 70.8 english⊗= 57.7 56.0 59.3 57.2 59.9 59.0 58.9 56.3 55.2 59.9 58.7 57.3 58.2 59.7 chinese⊗= 74.6 72.7 76.3 71.0 76.1 74.5 74.7 72.3 68.5 74.2 76.3 74.3 70.5 74.9
Type Dataset MCL XL Hindi AR EN FR RU ZH BG DA DE EN ET FA FR HR IT JA KO NL ZH HI mono hindi 56.8 56.6 50.9 50.1 49.2 50.3 52.7 49.5 56.3 51.5 50.5 51.3 50.5 57.4 52.7 53.0 53.5 48.8 86.5 english 56.2 86.6 64.5 50.7 49.1 50.2 52.4 52.5 70.4 51.0 49.1 56.0 46.3 53.4 50.0 50.0 51.0 50.1 81.4 french 52.7 61.2 56.7 50.3 50.0 52.8 56.1 48.8 58.7 54.9 51.5 50.9 50.7 54.2 50.8 59.6 53.8 54.0 64.0 german 55.3 68.5 62.7 51.0 51.2 49.9 56.0 80.0 62.0 55.1 56.5 57.4 48.8 55.2 51.1 50.0 56.4 51.0 69.8 chinese 50.0 50.2 52.8 50.9 54.3 49.9 51.4 51.6 51.7 52.1 51.9 51.6 53.2 55.4 50.6 58.4 51.5 52.1 56.6 multi multi 48.3 48.4 49.9 50.2 49.0 50.7 50.1 50.9 48.8 49.0 46.8 51.6 50.5 51.2 52.7 51.5 47.2 49.8 45.9
Type Dataset AM2iCO AR BN DE EU FI ID JA KA KK KO RU TR UR ZH mono hindi 49.6 50.3 50.8 49.5 52.1 48.5 47.5 49.9 49.0 50.8 50.6 50.2 50.7 50.0 english 53.7 71.7 65.3 55.9 57.3 59.8 49.8 49.0 49.0 49.3 51.7 54.8 70.2 48.7 french 49.5 51.1 51.6 48.8 49.7 49.9 50.0 48.1 50.7 47.5 51.9 51.4 51.5 50.2 german 52.6 57.6 56.9 54.8 54.4 56.2 51.1 52.2 51.8 50.2 53.7 52.5 53.8 51.1 chinese 50.0 50.1 50.0 50.0 49.9 50.0 50.0 50.0 50.0 50.0 49.9 50.0 50.0 50.0 multi multi 49.6 47.7 50.4 48.5 49.5 49.3 47.7 48.9 51.5 50.2 49.1 49.3 49.2 48.1
Appendix H Performance of XLM-R, mBERT and BLOOM-Based Models on LSCD Tasks
MuRIL models were excluded from this evaluation due to poor cross-lingual performance, resulting from a lack of pretraining on most of the target languages. LLaMA models were also not evaluated, as the LSCD task format is not well-suited to generative architectures.
| Language | Metric | xl-lexeme | multi | german | french | english | hindi | chinese |
| English | APD | .757 | .703 | .737 | .681 | .772 | .436 | .673 |
| PRT | .495 | .492 | .241 | .337 | .535 | .363 | .367 | |
| German | APD | .873 | .863 | .841 | .867 | .844 | .635 | .641 |
| PRT | .881 | .890 | .829 | .831 | .873 | .755 | .682 | |
| Swedish | APD | .755 | .801 | .754 | .618 | .724 | .480 | .489 |
| PRT | .678 | .673 | .522 | .138 | .627 | .277 | .332 | |
| Latin | APD | -.035 | .117 | .161 | .136 | .135 | -.177 | .091 |
| PRT | .467 | .392 | .445 | .405 | .429 | .512 | .512 | |
| Spanish | APD | .665 | .696 | .670 | .664 | .711 | .354 | .383 |
| PRT | .633 | .698 | .655 | .649 | .643 | .355 | .267 | |
| Chinese | APD | .734 | .652 | .649 | .499 | .737 | .524 | .593 |
| PRT | .702 | .708 | .623 | .578 | .684 | .552 | .432 | |
| Norwegian1 | APD | .668 | .729 | .638 | .697 | .777 | .525 | .400 |
| PRT | .769 | .784 | .730 | .740 | .845 | .551 | .435 | |
| Norwegian2 | APD | .634 | .655 | .604 | .580 | .645 | .433 | .439 |
| PRT | .532 | .583 | .557 | .525 | .636 | .337 | .396 | |
| Average | APD | .631 | .652 | .632 | .593 | .668 | .401 | .464 |
| PRT | .645 | .652 | .575 | .525 | .659 | .463 | .428 |
| Language | Metric | multi | german | french | english | hindi | chinese |
| English | APD | .754 | .566 | .551 | .711 | .506 | .684 |
| PRT | .470 | .397 | .393 | .434 | .309 | .328 | |
| German | APD | .760 | .719 | .772 | .810 | .551 | .783 |
| PRT | .834 | .759 | .846 | .845 | .550 | .705 | |
| Swedish | APD | .377 | .473 | .002 | .437 | .523 | .541 |
| PRT | .082 | .145 | -.504 | .266 | .207 | .366 | |
| Latin | APD | .181 | -.092 | -.058 | -.158 | .163 | -.190 |
| PRT | .298 | .339 | .235 | .038 | .462 | .412 | |
| Spanish | APD | .629 | .500 | .531 | .651 | .452 | .536 |
| PRT | .636 | .462 | .546 | .617 | .332 | .407 | |
| Chinese | APD | .654 | .554 | .386 | .668 | .651 | .619 |
| PRT | .690 | .623 | .493 | .686 | .427 | .529 | |
| Norwegian1 | APD | .596 | .515 | .561 | .638 | .631 | .700 |
| PRT | .724 | .698 | .580 | .748 | .558 | .601 | |
| Norwegian2 | APD | .601 | .596 | .464 | .604 | .412 | .581 |
| PRT | .474 | .527 | .260 | .535 | .238 | .421 | |
| Average | APD | .569 | .479 | .401 | .545 | .486 | .532 |
| PRT | .526 | .494 | .356 | .521 | .385 | .471 |
| Language | Metric | multi | german | french | english | hindi | chinese |
| English | APD | .689 | .463 | .657 | .656 | .176 | .291 |
| PRT | .469 | .293 | .459 | .487 | .215 | .342 | |
| German | APD | .634 | .606 | .244 | .410 | .075 | .144 |
| PRT | .720 | .723 | .474 | .564 | .274 | .499 | |
| Swedish | APD | .278 | .399 | .303 | .311 | .276 | .132 |
| PRT | .287 | .326 | .219 | -.052 | .134 | -.041 | |
| Latin | APD | -.068 | -.104 | .081 | -.197 | -.230 | -.173 |
| PRT | .367 | .221 | .323 | .314 | .125 | .310 | |
| Spanish | APD | .501 | .370 | .474 | .599 | .326 | .230 |
| PRT | .489 | .372 | .450 | .521 | .284 | .178 | |
| Chinese | APD | .550 | .467 | .570 | .592 | .216 | .299 |
| PRT | .546 | .381 | .506 | .568 | .395 | .159 | |
| Norwegian1 | APD | .278 | .289 | .131 | .345 | -.002 | .051 |
| PRT | .251 | .512 | -.015 | .354 | .240 | .363 | |
| Norwegian2 | APD | .395 | .193 | .237 | .089 | -.109 | -.046 |
| PRT | .264 | .216 | .261 | .041 | -.303 | -.118 | |
| Average | APD | .407 | .335 | .337 | .351 | .091 | .116 |
| PRT | .424 | .380 | .335 | .350 | .170 | .211 |
Appendix I LLaMA
I.1 LLaMA outperformed by XLM-R
Figure 7 shows that XLM-R generally outperforms LLaMA across most datasets under both full and fixed fine-tuning conditions, except when the models are trained on Hindi or Chinese. This can be attributed to LLaMA’s instruction-following pretraining, which enables it to better generalize from limited examples, which is particularly valuable in low-resource settings. In contrast, XLM-R lacks such capabilities and must infer both the task and the language from sparse training data. This issue is exacerbated by the fact that LLaMA was provided with prompts, which were consistently in English due to tokenizer limitations, while XLM-R had no explicit indication of either the task or input language during training or evaluation except for the sparse training data.
While we include LLaMA in our evaluation to provide a comparison with recent large language models and SOTA instruction-tuned architectures, our results suggest that encoder-based models like XLM-R remain more effective for in-context polysemy disambiguation. Despite the generative capabilities of LLaMA, it underperforms in most cases compared to XLM-R, particularly when ample training data is available. This highlights the continued relevance of encoder-based models, which demonstrate stronger task-specific performance and greater cross-lingual transfer.
I.2 Prompts and Response Parsing
System message:
Act as an expert lexicographer: determine if a given word has the same sense in two sentences, and respond with 1 if the sense is the same, or 0 if it is different.
User message:
Determine if "{target_word}" has the same meaning in these two sentences. You MUST reply with ONLY: 1 if the meaning is identical. 0 if the meaning differs. Sentence 1: "{sentence1}" Sentence 2: "{sentence2}" Answer (ONLY 1 or 0):
We adapted the WiC task to a generative setting for LLaMA by using the prompt shown in Figure 8. This prompt format was used consistently during both training and evaluation. For training, the prompt was followed by the correct binary label (“1” or “0”) as the target output. During evaluation, the model generated this label based on the prompt alone. We evaluated the pretrained model using several alternative prompt formulations on evaluation in all languages, ultimately selecting the one that yielded the highest overall performance.
To extract model predictions, we used the following simple rule-based parser:
if "1" in raw and "0" not in raw:
return "1"
elif "0" in raw and "1" not in raw:
return "0"
else:
return None # counts as incorrect
In practice, nearly all outputs followed the expected format, with the vast majority consisting of a single “1” or “0”.
Appendix J LLaMA Fine-Tuning Parameters
| Fine-tuning details | |||
| pre-trained LLMs |
| GPUs | NVIDIA A40 (48GB) | ||
| PEFT | LoRA | ||
| LoRA dropout | 0.1 | ||
| Weight decay | 0.001 | ||
| Learning rate | 1e-4 | ||
| Lora rank | 128 | ||
| Lora alpha | 256 | ||
| Warmup ratio | 0.05 | ||
| Num train epochs | 3 | ||
| Gradient accumulation steps | 4 | ||
| Max seq. length | 512 | ||
| Batch size | 8 | ||
| Optimizer | paged_adamw_8bit | ||
| LoRA target modules |